OpenStudy (anonymous):

2,4,6,8,16 next three numbers and rule

OpenStudy (anonymous):

n=2n

OpenStudy (anonymous):

32 64 128

OpenStudy (anonymous):

well actually its just 2n

OpenStudy (anonymous):

2, 2+2=4 ,4+2=6 ,6+2=8 , 8+4=16 ,16+2=18 , 18+2=20 20+2=22

OpenStudy (anonymous):

nope sorry just realised its not. Are you sure those are the correct numbers